# SketchOGD: Memory-Efficient Continual Learning

### v2

@@ IV-B Deriving Bounds
Sketching Method 4 is a modification of Sketching Method 3 to take advantage of the symmetric structure of $GG^{\top}$. It turns out that the metric for Sketching Method 4 is always better than for Sketching Method 3 as below, so we can transfer our previous results. See the appendix for the proof.
Theorem IV.4 (Error in Sketching Method 4).
When sketching $A=GG^{\top}$ into $\widehat{A}$ and $\widehat{A}_{\mathrm{\mathrm{sym}}}$ according to Sketching Methods 3 and 4, the metrics obey: $\mathcal{E}_{G}(\widehat{A}_{\mathrm{sym}})\leq\mathcal{E}_{G}(\widehat{A})$.
Consequently, the bounds described in Theorems IV.2 and IV.3 are also upper bounds for the error metric $\mathcal{E}_{G}(\widehat{A}_{\mathrm{sym}})$ and its expected value $\mathbb{E}[\mathcal{E}_{G}(\widehat{A}_{\mathrm{sym}})]$, respectively.
@@
Overall, these bounds suggest that the performance of different sketching methods may vary with the task-dependent structure of spectral decay.
V Experiments
In this section, we demonstrate the practical performance of SketchOGD on four continual learning benchmarks: Rotated MNIST, Permuted MNIST, and Split MNIST, and Split CIFAR.
@@ V Experiments
For the setup of our proposed SketchOGD methods, the memory constraints allow $k=1200$ for Method 1, $k=600$ for Method 2, and $k=299,l=301$ for Method 3. We set $k\approx l$ which is empirically found to be more memory-efficient. Effectively, SketchOGD-1 form a rank-$1200$ approximation, while SketchOGD-2 and SketchOGD-3 form around rank-$600$ approximations.
@@
For MNIST experiments, we use a three-layer fully connected neural network that has $100$ hidden units in the first two layers with ReLU activations and $10$ output logits, totaling $p=113,610$ parameters. For Split CIFAR, we use a LeNet with a hidden-layer size of $500$ for a total of $1,601,500$ parameters. Unlike previous works [4, 31] that create unique heads for each task in Split MNIST, we keep the model fixed to make Split MNIST a harder problem. However, we keep the multi-head architecture for Split CIFAR.
@@
We compare the SketchOGD methods with other continual learning methods under the same memory constraint. They include naive SGD for context, a variant of OGD which we call “RandomOGD” that randomly samples gradients to obey the memory constraint, PCA-OGD proposed by [31] that compresses each task’s gradients using PCA, and OGD without memory constraint. We expect naive SGD to perform drastically worse than all other methods because of catastrophic forgetting. RandomOGD should serve as a lower bound, as random sampling is the simplest form of compression. Unconstrained OGD serves as the theoretical maximum if compression methods were lossless, and PCA-OGD is the current state-of-the-art memory-efficient OGD variant.
@@
V-A Experiment 1: All Methods under the Same Memory Constraint
@@
V-B Experiment 2: Compressing the Same Number of Gradients
@@ VI Discussion
Our results motivate several directions of future inquiry. First, Theorem IV.4 leaves room for a more rigorous analysis of the benefits of leveraging the symmetric structure of $GG^{\top}$ in SketchOGD-3. Also, the difference between the bounds in IV.1 and IV.3 involve replacing one instance of $\operatorname{Tr}(\Sigma_{2})$ with $\operatorname{Tr}(\Sigma_{2}^{2})\operatorname{Tr}(\Sigma_{1}^{-1})$; we would like to investigate the way different types of singular value decay would affect these two bounds differently. Another area of importance is one of our potential explanations for SketchOGD-2,3’s performance, namely the quadratically scaled magnitudes of gradients in $GG^{\top}$. We may test whether scaling the gradients by powers of their magnitudes before compression may influence results for the better. In addition, given SketchOGD’s incremental nature, developing an incremental PCA-OGD variant for comparison is another promising direction However, any incremental forms of PCA won’t have the linear properties of sketching that make it equivalent to sketching the full matrix all at once; a modified PCA will likely still involve repeated truncation, discarding important directions before they have time to accumulate across tasks.
VII Conclusion
@@ Acknowledgements
This work was supported in part by the MIT-IBM Watson AI Lab, MathWorks, the MIT-Amazon Science Hub, and the MIT-Google Program for Computing Innovation. The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center [33] for providing computing resources that have contributed to the research results reported within this paper.
@@ A-A1 Effects of Spectral Decay on Bounds
Plugging $\gamma^{*}$ in, we get the upper bound $(1+\tfrac{\gamma^{*}}{k-\gamma^{*}-1})\cdot\operatorname{Tr}{\Sigma_{2}}=p\lambda$. Intuitively, a flat decay pushes the optimal split location $\gamma$ to 00, and scaling the singular values by a constant does not affect the optimal split location.
In contrast, let’s consider $\Sigma$ of which the diagonal values linearly decay from $2\lambda$ to 00 over $p$ values. Then, the optimal split location can change:
@@
where the approximation for calculating $\operatorname{Tr}{\Sigma_{2}}$ comes from the simplification $21+2+\ldots+(p-\gamma)=(p-\gamma)^{2}/$ 2.
To calculate when the minimum is achieved, we can set the derivative with respect to $\gamma$ to be 00:
@@
which results in $\gamma^{*}=2k-2-p$. As long as $2k-2-p\geq 0$, then the optimal split location then becomes $\gamma^{*}=2k-2-p$. Plugging in, the upper bound computes to $\frac{4\lambda}{p}(k{-}1)(p{-}k{+}1)$. Note that this bound is smaller than in the constant singular values case, even though the sum of the singular values is the same. Similarly, the faster the decay becomes, the more $\gamma^{*}$ is pushed away from 00, and the tighter the theoretical upper bound becomes.
@@ A-B Proofs
$\mathbb{E}[\mathcal{E}_{G}(\widehat{G})]=\mathbb{E}\|(G-QQ^{\top}G)\|_{F}^{2}.$
Then, we can apply Proposition 10.5 of Halko et al. [28]:
$\mathbb{E}[\mathcal{E}_{G}(\widehat{G})]\leq\min_{\gamma<k-1}(1+\tfrac{\gamma}{k-\gamma-1})\cdot\tau^{2}_{\gamma+1}G,$ (1)
@@
By unitary invariance of the Frobenius norm we can drop the leading orthogonal factors: $\widetilde{A}=\Sigma U^{\top}=U^{\top}A$, $\widetilde{Y}=\widetilde{A}\Omega$, $\widetilde{G}=U^{\top}G=\Sigma^{1/2}V$.
$\|(1-\operatorname{proj}_{\widehat{A}})$ $G\|_{F}^{2}=\|(1-\operatorname{proj}_{Y})G\|_{F}^{2}$
@@
$\operatorname{range}\begin{bmatrix}\Sigma_{1}U_{1}^{\top}\\ 0\end{bmatrix}=\operatorname{range}\begin{bmatrix}\Sigma_{1}\Omega_{1}\\ 0\end{bmatrix}=\operatorname{range}(\widetilde{Y}),$
@@
where $F=\Sigma_{2}\Omega_{2}\Omega_{1}^{\dagger}\Sigma_{1}^{-1}$. By construction, $\operatorname{range}(Z)\subseteq\operatorname{range}(\widetilde{Y})$, so $Z^{\perp}$ can be split into $Y^{\perp}$ and an orthogonal subspace: $Z^{\perp}=Y^{\perp}\oplus(Y\cap Z^{\perp})$. Therefore, the projection of any vector onto $Z^{\perp}$ is at least as large as the projection onto $Y^{\perp}$. Apply this result to every column of $\widetilde{G}$ to get
@@
By Proposition 8.2 of Halko et al. [28], $I-(I+F^{\top}F)^{-1}\preccurlyeq F^{\top}F$. Next, as $F^{\top}F$ is PSD and symmetric, $(I+F^{\top}F)^{-1}$ is also PSD and symmetric. So, by the conjugation rule, $F(I+F^{\top}F)^{-1}F^{\top}$ is PSD. Therefore $I-F(I+F^{\top}F)^{-1}F^{\top}\preccurlyeq I$. So,
$1-\operatorname{proj}_{{Z}}\preccurlyeq\begin{bmatrix}F^{\top}F&B\\ B^{\top}&I\end{bmatrix}.$
@@
$\Sigma^{1/2}(1-\operatorname{proj}_{{Z}})\Sigma^{1/2}\preccurlyeq\begin{bmatrix}\Sigma^{1/2}_{1}F^{\top}F\Sigma^{1/2}_{1}&\Sigma^{1/2}_{1}B\Sigma^{1/2}_{2}\\ \Sigma^{1/2}_{2}B^{\top}\Sigma^{1/2}_{1}&\Sigma^{1/2}_{2}\Sigma^{1/2}_{2}\end{bmatrix}.$
@@
$\mathcal{E}_{G}(\widehat{A})$ $\leq\|F\Sigma^{1/2}_{1}\|^{2}_{F}+\|\Sigma^{1/2}_{2}\|^{2}_{F}$
@@
$\mathcal{E}_{G}(\widehat{A})\leq\|\Sigma_{2}\Omega_{2}\Omega_{1}^{\dagger}\Sigma_{1}^{-1/2}\|^{2}+\|\Sigma_{2}^{1/2}\|^{2},$
@@
$\mathbb{E}[\mathcal{E}_{G}(\widehat{A})]\leq\mathbb{E}\left[\|\Sigma_{2}\Omega_{2}\Omega_{1}^{\dagger}\Sigma_{1}^{-1/2}\|^{2}+\|\Sigma_{2}^{1/2}\|^{2}\right].$
The expectation can be split over $\Omega_{2}$ and $\Omega_{1}$. Now, we use Proposition 10.1 of Halko et al. [28], which states that $\mathbb{E}_{G}\|AGB\|^{2}=\|A\|^{2}\|B\|^{2}$ if $G$ is standard normally distributed:
@@
$=\mathbb{E}_{\Omega_{1}}\left[\|\Sigma_{2}\|^{2}\|\Omega_{1}^{\dagger}\Sigma_{1}^{-1/2}\|^{2}+\|\Sigma_{2}^{1/2}\|^{2}\right]$
$=\operatorname{Tr}(\Sigma_{2}^{2})\mathbb{E}_{\Omega_{1}}\left[\|\Omega_{1}^{\dagger}\Sigma_{1}^{-1/2}\|^{2}\right]+\operatorname{Tr}(\Sigma_{2}).$
@@
Now, note that $W=\Omega_{1}\Omega_{1}^{\top}$ has a Wishart distribution, namely $W_{k}(\gamma,I)$, so, by a well-known result of multivariate statistics [34, pg. 97], $\mathbb{E}[W^{-1}]=\frac{1}{k-\gamma-1}I$. Then:
@@
$\mathbb{E}[\mathcal{E}_{G}(\widehat{A})]\leq\min_{\gamma\leq k-2}\tfrac{\gamma}{k-\gamma-1}\operatorname{Tr}(\Sigma_{2}^{2})\operatorname{Tr}(\Sigma_{1}^{-1})+\operatorname{Tr}(\Sigma_{2}).$
@@
For matrices $Q,X$ described in Definition 3.3, $\operatorname{range}(\widehat{A}_{sym})=\operatorname{range}([Q,X^{\top}])=\operatorname{range}(\widehat{A})+\operatorname{range}(X^{\top})$. So, the range of $\widehat{A}$ is just a subset of the range of the symmetric approximation $\widehat{A}_{sym}$, which implies:
$|(1-\operatorname{proj}_{\widehat{A}_{sym}})u|_{2}\leq|(1-\operatorname{proj}_{\widehat{A}})u|_{2}.$

### v3

@@ IV-B Deriving Bounds
Sketching Method 3 is a modification of Sketching Method 2 to take advantage of the symmetric structure of $GG^{\top}$ via the symmetric sketch in Proposition II.3. Sketching Method 3 returns a basis that spans a larger subspace than that from Sketching Method 2. Any projection onto this larger subspace can only reduce the reconstruction error compared to Sketching Method 2. Thus, the metric for Sketching Method 3 is always better than for Sketching Method 2 as below, so we can transfer our previous results. See the appendix for the proof.
Theorem IV.4 (Error in Sketching Method 3).
When sketching $A=GG^{\top}$ to obtain orthonormal bases $B$ and $B_{\mathrm{sym}}$ according to Sketching Methods 2 and 3, respectively, the metrics obey:
$\mathcal{E}_{G}(B_{\mathrm{sym}})\leq\mathcal{E}_{G}(B).$
Consequently, the bounds described in Theorems IV.2 and IV.3 are also upper bounds for the error metric $\mathcal{E}_{G}(B_{\mathrm{sym}})$ and its expected value $\mathbb{E}[\mathcal{E}_{G}(B_{\mathrm{sym}})]$, respectively.
@@
Overall, these bounds suggest that the performance of different sketching methods may vary with the task-dependent structure of spectral decay. More sharply decaying spectrum of $GG^{\top}$ makes the bound for Sketching Methods 2 and 3 in Theorem IV.3 tighter than that for Sketching Method 1 in Theorem IV.1. Conversely, when the spectrum decays slowly, Sketching Method 1 can be more favorable. Among Sketching Methods 2 and 3, the metric for Sketching Method 3 is always better than Sketching Method 2 as noted in Theorem IV.4.
V Analysis of Convergence and Catastrophic Forgetting
In this section, we theoretically characterize the limiting parameter attained by SketchOGD on each task and analyze the catastrophic forgetting induced by the successive parameter updates along tasks.
Under the Neural Tangent Kernel (NTK) regime, bennani2020generalisation and Doan2020ATA interpret OGD as a recursive kernel regression and analyze its forgetting behavior. We adopt a similar approach and establish an explicit connection between the sketching reconstruction error $\mathcal{E}_{G}(B)$ and the catastrophic forgetting.
The key idea of the NTK regime is that, for a neural network with large enough width, the model is well-approximated by its first-order approximation around the initialization $w_{0}\in\mathbb{R}^{p}$:
$f(x,w)\approx f(x,w_{0})+\nabla_{w}f(x,w_{0})^{\top}(w-w_{0}),$ (1)
where the feature map $\nabla_{w}f(x,w)$ remains constant throughout training [JacotNTK, LeeNTK]. This yields a kernel regression viewpoint of learning and enables closed-form characterizations of the solutions reached by SketchOGD.
For each task $\tau\in\{1,2,\dots,T\}$, we train on the dataset $\{(x_{\tau,i},y_{\tau,i})\}_{i=1}^{n_{\tau}}$, arranged as $\mathtt{X}_{\tau}\in\mathbb{R}^{d\times n_{\tau}}$, and $\mathtt{Y}_{\tau}\in\mathbb{R}^{n_{\tau}}$, with regularized squared loss function:
$\mathcal{L}_{\tau}(w)=\frac{1}{n_{\tau}}\|f(\mathtt{X}_{\tau},w)-\mathtt{Y}_{\tau}\|_{2}^{2}+\lambda\|w-w_{\tau-1}^{*}\|_{2}^{2},$ (2)
where $f(\mathtt{X}_{\tau},w)\in\mathbb{R}^{n_{\tau}}$ is an aggregated function evaluation on each input of $\mathtt{X}_{\tau}$. Let $G_{\tau}\in\mathbb{R}^{p\times n_{\tau}}$ denote the matrix whose columns are the gradients for the $n_{\tau}$ training data. Prior to training task $\tau$, a sketching method constructs $B_{\tau}$ based on the gradients of prior tasks $G_{1:\tau-1}$. Denoting the orthogonal projection as $T_{\tau}:=I_{p}-B_{\tau}B_{\tau}^{\top}$, each $j$-th update step of SketchOGD can be represented as:
$w_{\tau}^{(j+1)}=w_{\tau}^{(j)}-\eta_{\tau}T_{\tau}\nabla_{w}\mathcal{L}_{\tau}(w_{\tau}^{(j)}),$ (3)
for a fixed step-size $\eta_{\tau}>0$.
Under this setting, we characterize the convergence behavior of SketchOGD as below. See the appendix for the proof.
Theorem V.1 (Convergence analysis).
Suppose the NTK assumption (1) holds and the step-size $\eta_{\tau}$ is sufficiently small so that $0<\eta_{\tau}<\dfrac{1}{\lambda_{max}\big(T_{\tau}(\frac{1}{n_{\tau}}G_{\tau}G_{\tau}^{\top}+\lambda I_{p})\big)}$ for all $\tau\in\{1,\dots,T\}$. Then, for each task $\tau$, the parameter updates $\{w_{\tau}^{(j)}\}_{j\geq 0}$ of SketchOGD in (3) converge linearly to a limit solution $w_{\tau}^{*}$ such that
$w_{\tau}^{*}-w_{\tau-1}^{*}=T_{\tau}G_{\tau}(G_{\tau}^{\top}T_{\tau}G_{\tau}+n_{\tau}\lambda I_{n_{\tau}})^{-1}\tilde{\mathtt{Y}}_{\tau},$ (4)
where $\tilde{\mathtt{Y}}_{\tau}:=\mathtt{Y}_{\tau}-f(\mathtt{X}_{\tau},w_{\tau-1}^{*})$ and $w_{0}^{*}=w_{0}$.
The recursive characterization of the limiting parameters in (4) enables a direct analysis of catastrophic forgetting. Similarly to the definition from Doan2020ATA, we define catastrophic forgetting on the training data as follows.
Definition V.1 (Catastrophic forgetting).
Let $\tau_{S}$ and $\tau_{T}$ ($\tau_{S}<\tau_{T}$) be the source and target tasks, respectively. The catastrophic forgetting of task $\tau_{S}$ after training on all subsequent tasks up to task $\tau_{T}$ is
$CF_{\tau_{S}\rightarrow\tau_{T}}=\sum_{i=1}^{n_{\tau_{S}}}\big(f(x_{\tau_{S},i},w_{\tau_{T}}^{*})-f(x_{\tau_{S},i},w_{\tau_{S}}^{*})\big)^{2},$
where $w_{\tau_{S}}^{*},w_{\tau_{T}}^{*}$ are the limiting parameters from Theorem V.1.
Then, with the NTK approximation, the catastrophic forgetting admits the equivalent form:
$CF_{\tau_{S}\rightarrow\tau_{T}}$ $=\|G_{\tau_{S}}^{\top}(w_{\tau_{T}}^{*}-w_{\tau_{S}}^{*})\|^{2}$
$=\big\|\sum_{\tau=\tau_{S}+1}^{\tau_{T}}G_{\tau_{S}}^{\top}(w_{\tau}^{*}-w_{\tau-1}^{*})\big\|^{2}.$
Combining this identity with the recursive characterization of $w_{\tau}^{*}$ in (4) yields the following results.
Theorem V.2 (Catastrophic forgetting of SketchOGD).
Suppose the conditions of Theorem V.1 hold. Then, for the source task $\tau_{S}$ and target task $\tau_{T}$ ($\tau_{S}<\tau_{T}$),
$CF_{\tau_{S}\rightarrow\tau_{T}}=\big\|\sum_{\tau=\tau_{S}+1}^{\tau_{T}}G_{\tau_{S}}^{\top}T_{\tau}G_{\tau}(G_{\tau}^{\top}T_{\tau}G_{\tau}+n_{\tau}\lambda I)^{-1}\tilde{\mathtt{Y}}_{\tau}\big\|^{2}.$
Remark V.2.
For each task $\tau>\tau_{S}$, the reconstruction error defined in Definition IV.1 upper bounds the magnitude of the cross-term $G_{\tau_{S}}^{\top}T_{\tau}$ since the error metric after task $\tau-1$ is
$\mathcal{E}_{G_{1:\tau-1}}(B_{\tau})=\|T_{\tau}G_{1:\tau-1}\|_{F}^{2}\geq\|G_{\tau_{S}}^{\top}T_{\tau}\|_{F}^{2}.$
A small reconstruction error implies that $G_{\tau_{S}}$ is nearly contained in the sketched subspace, thereby leading to smaller catastrophic forgetting.
VI Experiments
In this section, we demonstrate the practical performance of SketchOGD on four continual learning benchmarks: Rotated MNIST, Permuted MNIST, Split MNIST, and Split CIFAR.
@@ VI Experiments
Our primary goal is to compare methods at fixed memory budgets, which is the typical design constraint in continual learning scenarios. For the setup of our proposed SketchOGD methods, the memory constraints allow $k=1200$ for Method 1, $k=600$ for Method 2, and $k=299,l=301$ for Method 3. We set $k\approx l$, which is empirically found to be more memory-efficient. Effectively, SketchOGD-1 forms a rank-$1200$ approximation, while SketchOGD-2 and SketchOGD-3 form around rank-$600$ approximations.
@@
For MNIST experiments, we use a three-layer fully connected neural network that has $100$ hidden units in the first two layers with ReLU activations and $10$ output logits, totaling $p=113,610$ parameters. For Split CIFAR, we use a LeNet with a hidden-layer size of $500$ for a total of $1,601,500$ parameters. Unlike previous works [Farajtabar2019OrthogonalGD, Doan2020ATA] that create unique heads for each task in Split MNIST, we keep the model fixed to make Split MNIST a harder problem. However, we keep the multi-head architecture for Split CIFAR.
@@
We compare the SketchOGD methods with other continual learning methods under the same memory constraint. They include naive SGD for context, a variant of OGD which we call “RandomOGD” that randomly samples gradients to obey the memory constraint, PCA-OGD proposed by [Doan2020ATA] that compresses each task’s gradients using PCA, and OGD without memory constraint. We expect naive SGD to perform drastically worse than all other methods because of catastrophic forgetting. RandomOGD should serve as a lower bound, as random sampling is the simplest form of compression. Unconstrained OGD serves as the theoretical maximum if compression methods were lossless, and PCA-OGD is the current state-of-the-art memory-efficient OGD variant.
@@
VI-A Experiment 1: All Methods under the Same Memory Constraint
@@
VI-B Experiment 2: Compressing the Same Number of Gradients
@@ VII Discussion
Our results motivate several directions of future inquiry. First, Theorem IV.4 leaves room for a more rigorous analysis of the benefits of leveraging the symmetric structure of $GG^{\top}$ in SketchOGD-3. Also, the difference between the bounds in IV.1 and IV.3 involve replacing one instance of $\operatorname{Tr}(\Sigma_{2})$ with $\operatorname{Tr}(\Sigma_{2}^{2})\operatorname{Tr}(\Sigma_{1}^{-1})$; we would like to investigate the way different types of singular value decay would affect these two bounds differently. Another area of importance is one of our potential explanations for SketchOGD-2,3’s performance, namely the quadratically scaled magnitudes of gradients in $GG^{\top}$. We may test whether scaling the gradients by powers of their magnitudes before compression may influence results for the better. In addition, given SketchOGD’s incremental nature, developing an incremental PCA-OGD variant for comparison is another promising direction. However, any incremental forms of PCA won’t have the linear properties of sketching that make it equivalent to sketching the full matrix all at once; a modified PCA will likely still involve repeated truncation, discarding important directions before they have time to accumulate across tasks. Another interesting direction is to analyze formal privacy guarantees [zhu2021rgap, wang2022protect], given that one of SketchOGD’s primary advantages is its ability to avoid storing raw data, which is a key source of privacy concerns. In addition, extending the empirical study to larger-scale datasets (e.g., CORe50, ImageNet) and modern architectures (e.g., ResNets, Transformers) represents an important avenue for future research.
VIII Conclusion
@@ Acknowledgements
This work was supported in part by the MIT-IBM Watson AI Lab, MathWorks, the MIT-Amazon Science Hub, and the MIT-Google Program for Computing Innovation. The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center [reuther2018interactive] for providing computing resources that have contributed to the research results reported within this paper.
@@ A-A1 Effects of Spectral Decay on Bounds
Plugging $\gamma^{*}$ in, we get the upper bound $(1+\tfrac{\gamma^{*}}{k-\gamma^{*}-1})\cdot\operatorname{Tr}{\Sigma_{2}}=p\lambda$. Intuitively, a flat decay pushes the optimal split location $\gamma$ to 0, and scaling the singular values by a constant does not affect the optimal split location.
In contrast, let’s consider $\Sigma$ of which the diagonal values linearly decay from $2\lambda$ to 0 over $p$ values. Then, the optimal split location can change:
@@
where the approximation for calculating $\operatorname{Tr}{\Sigma_{2}}$ comes from the simplification $1+2+\ldots+(p-\gamma)=(p-\gamma)^{2}/2$.
To calculate when the minimum is achieved, we can set the derivative with respect to $\gamma$ to be 0:
@@
which results in $\gamma^{*}=2k-2-p$. As long as $2k-2-p\geq 0$, then the optimal split location becomes $\gamma^{*}=2k-2-p$. Plugging in, the upper bound computes to $\frac{4\lambda}{p}(k{-}1)(p{-}k{+}1)$. Note that this bound is smaller than in the constant singular values case, even though the sum of the singular values is the same. Similarly, the faster the decay becomes, the more $\gamma^{*}$ is pushed away from 0, and the tighter the theoretical upper bound becomes.
@@ A-B Proofs
$\mathbb{E}[\mathcal{E}_{G}(B)]=\mathbb{E}\|(G-BB^{\top}G)\|_{F}^{2}.$
Then, we can apply Theorem 10.5 of halko:
$\mathbb{E}[\mathcal{E}_{G}(B)]\leq\min_{\gamma<k-1}(1+\tfrac{\gamma}{k-\gamma-1})\cdot\tau^{2}_{\gamma+1}G,$ (5)
@@
By unitary invariance of the Frobenius norm, we can drop the leading orthogonal factors: $\widetilde{A}=\Sigma U^{\top}=U^{\top}A$, $\widetilde{Y}=\widetilde{A}\Omega$, $\widetilde{G}=U^{\top}G=\Sigma^{1/2}V$.
$\|(1-\operatorname{proj}_{B})$ $G\|_{F}^{2}=\|(1-\operatorname{proj}_{Y})G\|_{F}^{2}$
@@
[…]
@@
[…]
@@
By Proposition 8.2 of halko, $I-(I+F^{\top}F)^{-1}\preccurlyeq F^{\top}F$. Next, as $F^{\top}F$ is PSD and symmetric, $(I+F^{\top}F)^{-1}$ is also PSD and symmetric. So, by the conjugation rule, $F(I+F^{\top}F)^{-1}F^{\top}$ is PSD. Therefore $I-F(I+F^{\top}F)^{-1}F^{\top}\preccurlyeq I$. So,
$1-\operatorname{proj}_{{Z}}\preccurlyeq\begin{bmatrix}F^{\top}F&M\\ M^{\top}&I\end{bmatrix}$.
@@
$\Sigma^{1/2}(1-\operatorname{proj}_{{Z}})\Sigma^{1/2}\preccurlyeq\begin{bmatrix}\Sigma^{1/2}_{1}F^{\top}F\Sigma^{1/2}_{1}&\Sigma^{1/2}_{1}M\Sigma^{1/2}_{2}\\ \Sigma^{1/2}_{2}M^{\top}\Sigma^{1/2}_{1}&\Sigma^{1/2}_{2}\Sigma^{1/2}_{2}\end{bmatrix}$.
@@
$\mathcal{E}_{G}(B)$ $\leq\|F\Sigma^{1/2}_{1}\|^{2}_{F}+\|\Sigma^{1/2}_{2}\|^{2}_{F}$
@@
$\mathcal{E}_{G}(B)\leq\|\Sigma_{2}\Omega_{2}\Omega_{1}^{\dagger}\Sigma_{1}^{-1/2}\|^{2}+\|\Sigma_{2}^{1/2}\|^{2},$
@@
$\mathbb{E}[\mathcal{E}_{G}(B)]\leq\mathbb{E}\left[\|\Sigma_{2}\Omega_{2}\Omega_{1}^{\dagger}\Sigma_{1}^{-1/2}\|^{2}+\|\Sigma_{2}^{1/2}\|^{2}\right].$
The expectation can be split over $\Omega_{2}$ and $\Omega_{1}$. Now, we use Proposition 10.1 of halko, which states that $\mathbb{E}_{G}\|M_{1}M_{2}M_{3}\|^{2}=\|M_{1}\|^{2}\|M_{3}\|^{2}$ if $M_{2}$ is standard normally distributed:
@@
$\hskip-56.9055pt=\mathbb{E}_{\Omega_{1}}\left[\|\Sigma_{2}\|^{2}\|\Omega_{1}^{\dagger}\Sigma_{1}^{-1/2}\|^{2}+\|\Sigma_{2}^{1/2}\|^{2}\right]$
$\hskip-56.9055pt=\operatorname{Tr}(\Sigma_{2}^{2})\mathbb{E}_{\Omega_{1}}\left[\|\Omega_{1}^{\dagger}\Sigma_{1}^{-1/2}\|^{2}\right]+\operatorname{Tr}(\Sigma_{2}).$
@@
Now, note that $W=\Omega_{1}\Omega_{1}^{\top}$ has a Wishart distribution, namely $W_{k}(\gamma,I)$, so, by a well-known result of multivariate statistics [muirhead, pg. 97], $\mathbb{E}[W^{-1}]=\frac{1}{k-\gamma-1}I$. Then:
@@
$\mathbb{E}[\mathcal{E}_{G}(B)]\leq\min_{\gamma\leq k-2}\tfrac{\gamma}{k-\gamma-1}\operatorname{Tr}(\Sigma_{2}^{2})\operatorname{Tr}(\Sigma_{1}^{-1})+\operatorname{Tr}(\Sigma_{2}).$
@@
For matrices $Q,X$ described in Sketching Method 3, $\operatorname{range}(B_{\mathrm{sym}})=\operatorname{range}([Q,X^{\top}])=\operatorname{range}(B)+\operatorname{range}(X^{\top})$. So, the range of $B$ is just a subset of the range of the symmetric approximation $B_{\mathrm{sym}}$, which implies:
$|(1-\operatorname{proj}_{B_{\mathrm{sym}}})u|_{2}\leq|(1-\operatorname{proj}_{B})u|_{2}.$